Fermi National Laboratory
 

Accelerator Update for June 29, 2007 - July 2, 2007

 

Friday June 29

The day shift began with Operations monitoring the Tevatron (TeV) store 5526 (93.77E30) and the Recycle stash (114.4E10), with the Antiproton Source (Pbar) stacking (10.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and the Meson MTest area taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 11:24 AM, RF experts bypassed a Booster RF station (BRF5) for repairs. They completed their repairs and brought the station back online at 1:30 PM.

At 2:53 PM, BRF5 tripped off. It reset.

Operations reported at 3:56 PM that the Recycler had stashed antiprotons twice during the day shift. Operations also reported that the TeV sector D1 wet engine had its connecting rod bearing fail. Cryo system technicians will replace the bearing in between stores.

MTest stopped taking beam for the night at 6:02 PM.

Operations terminated store 5526 at 8:01 PM.

Cryo system techs began their work on the D1 wet engine at 8:08 PM. The repair took a little over two hours to complete.

Operations prepared D0 for access at 8:22 PM.

At 10:38 PM, BRF13 tripped off on a ground fault. It reset.

At 10:47 PM, BRF13 tripped off. Operators bypassed the station and brought BRF19 online in it place.

Operations put the TeV through a dry squeeze at 11:45 PM.

Operations reported at 11:59 PM that the Recycler had stashed twice during the evening shift.

Operations put the TeV into shot setup at 11:59 PM.

Saturday June 30

The midnight shift began with Operations monitoring the stash (297.3E10), with Pbar stacking (37.9E10), and with NuMI and MiniBooNE taking beam.

At 12:09 AM, at Linac Klystron RF station (KRF2) tripped off and would not initially reset, but eventually did.

At 12:29 AM, the Pbar Lithium lens tripped off. Operations contacted experts.

A Pbar expert had the lens up and running by 1:33 AM. Stacking resumed.

Operations established store 5528 at 2:27 AM with an initial luminosity of 206.3E30.

Operations established beam to the MTest experiment at 6:25 AM.

Operations reported at 8:10 AM that the Recycler had stashed three times during the midnight shift.

At 10:34 AM, the Linac 750KeV buncher tripped off. It reset.

At 2:42 PM, Operations reported that they had been suffering Pbar control problems for about an hour. Operations contacted an expert who had the problem fixed by 3:17 PM.

Operations disabled the MTest beam at 6:22 PM.

Operations reported at 7:55 PM that the Recycler had stashed four times during the last 12 hours. Accelerator status: store 5528 was at 47.28E30, the stack was at 37.5E10, and the stash was at 273.0E10.

Operations terminated store 5528 at 8:59 PM.

Operations prepared D0 for access at 9:07 PM.

At 9:24 PM, Operations reported problems with the H- Source. Operators contacted an expert and then replaced its control card, which fixed the problem.

At 9:45 PM, BRF6 tripped off. It reset.

Operations put the TeV through a dry squeeze at 10:07 PM.

Operations put the TeV into shot setup at 10:35 PM.

Sunday July 1

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations established store 5530 at 12:38 AM with an initial luminosity of 204.78E30.

At 1:46 AM, the MI North Anode Power Supply tripped off and took down all the MI RF stations. Operators rebooted the Low Level RF, but it didn’t take. Operations contacted an expert who instructed them on the process to get everything running, which Operations followed and had everything running by 2:20 AM.

Operations established beam to MTest at 6:30 AM.

Operations reported at 7:55 AM that the Recycler had stashed five times during the last 12 hours. Accelerator status: store 5530 was at 82.8E30, the stack was at 12.0E10, and the stash was at 151.7E10.

At 8:55 AM, Switchyard (SY) tripped off due to a vacuum trip (FSEPV). It reset.

At 12:14 PM, Operations reported that the Recycler gating was found on (it’s supposed to remain off) for the first of three times over the next eight hours. It was shut off each time.

At 5:28 PM, a Meson quadrupole power supply (MT5Q2) tripped off due to magnet Klixon interlock. It would not reset.

Operations disabled the MTest beam at 6:33 PM.

The MT5Q2 magnet cooled off enough by 6:57 PM to reset and turn on.

Operations reported at 7:58 PM that the Recycler had stashed four times during the last 12 hours.

Operations terminated store 5530 at 8:02 PM.

Operations put the TeV into shot setup at 8:15 PM.

At 8:31 PM, Operations reported a problem with the TeV sector F4 wet engine. An expert was called in to investigate. Shot setup was put on hold.

The Cryo Coordinator reported at 9:54 PM that Operations could continue with shot setup.

Operations established store 5531 at 11:34 PM with an initial luminosity of 201.15E30.

Monday July 2

The midnight shift began with Operations monitoring store 5531 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 1:04 AM, Operations reported that the Linac buncher had tripped off. It reset.

At 1:20 AM, the TeV sector C17 separator sparked two times.

Operations prepared MTest for beam at 6:12 AM.

Operations reported at 7:57 AM that the Recycler had stashed four times during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Linac, Pbar, and Booster experts to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/02/2007    email Fermilab